LHC 2010 - latest news

Week 21: Machine Coordinators:    Mike Lamont - Ralph Assmann

Latest plan: link PDF link XLS (version of 27 May, 00:30)

List of End of Fill studies (in work)

List of MD requests (in work)

Day Short summary/plan Link to reports from 8:30 and 17:00 meetings
Sunday 30th May
  • Recovery from power cut.
  • Aim for beam tests. No physics.
Report from 9h00 meeting
Saturday 29th May
  • All day: Recovery from power cut
Report from 9h00 meeting
Friday 28th May
  • 04h00: Study of BPM shifts with intensity.
  • 09h30: Access for QPS.
  • 11h30: pre-cycle.
  • 13h00: Impedance study with collimation.
  • 18h00: Test ramp of 1e11 bunch to 3.5 TeV. Pilot collsions.
  • 23h00: Rampdown and recover.
  • 23h20: General power cut.
Report from 8h30 meeting
Thursday 27th May
  • 00h15: Starting high intensity injections.
  • 00h34: Sector 67 tripped again (see 21h51 entry Wednesday). Access. BLM problem.
  • 08h52: Power glitch affecting various systems: RF, collimators, LHCb and ALICE dipoles, SPS. Recovery.
  • 13h24: Injection setup. Injection 7 x 1e11 per beam.
  • 14h59: Stable beams (450 GeV collisions).
  • 16h03: End of fill due to beam excursion interlock. Recovery. Sector 12 trip.
  • 17h40: Sector 67 trip.
  • 19h00: Beam back.
  • 20h38: Transverse damper studies.
Report from 8h30 meeting
Wednesday 26th May
  • 00h50: Collapsing separation. Checking collisions with nominal bunches. Large losses.
  • 03h30: Requested beam dump. Recover.
  • 04h30: Several trips and kicker MKI.UA87 problem. Access for power converter RD2.R8.
  • 10h20: Precycle and set up injection.
  • 15h00: Start ramp of 1 nominal bunch per beam. Beam 2 dumped during ramp.
  • 15h56: Requested dump of beam 1. Rampdown and recovery.
  • 18h53: Start ramp of 1 nominal bunch per beam.
  • 19h25: Arrive at 3.5 TeV with two beams. Adjust chromaticity and tunes (collision tunes).
  • 20h41: Collisions in CMS. Optimization.
  • 21h04: Collisions in CMS and ATLAS. Optimization.
  • 21h38: End of fill RF studies.
  • 21h51: Beam dump (18 kV glitch). Recovery. Ramp down (RB trip S67). Injection set up.
  • 23h56: Injection of pilot beams.
Report from 8h30 meeting
Tuesday 25th May
  • 01h11: Start ramp with 13 bunches, each 2e10
  • 03h16: Stable beams (13 x 2e10 p per beam)
  • 12h30: Requested beam dump.
  • 15h00: Energy matching beam 1.
  • 17h00: Injection kicker studies beam 2.
  • 18h00: Nominal bunch studies: RF
  • 23h30: Ramp nominal bunch to 3.5 TeV.
Report from 8h30 meeting
Monday 24th May
  • 00h30: Beam lost - LBDS dump.
  • 03h00: Injection. Again LBDS dump.
  • 04h00: Access for LBDS
  • 09h00: Injection.
  • 12h30: Start ramp with 13 bunches, each 2e10
  • 15h00: Stable beams (13 x 2e10 p per beam)
  • 17h30: Beam dump due to BLM card problem. Recover. Work on BLM cards. Injection setup.
Report from 9h00 meeting